OpenStudy (anonymous):

If h(x)=(fog)(x) and h(x)=5(x+1)^3, find f(x) and g(x). A. f(x)=5x^3, g(x)=(x+1)^3 B. f(x)=5x^3, g(x)=x+1 C. f(x)=x+1, g(x)=5x^3 D. f(x)=(x+1)^3, g(x)=5x^3

ganeshie8 (ganeshie8):

find (fog)(x) for each options and see which one gives u 5(x+1)^3

ganeshie8 (ganeshie8):

A. f(x)=5x^3, g(x)=(x+1)^3 (fog)(x) = f(g(x)) = f((x+1)^3) = 5(x+1)^6 REJECT

ganeshie8 (ganeshie8):

similarly test other options and see :)

OpenStudy (anonymous):

how do i "fog" the other equations though? :o @ganeshie8

ganeshie8 (ganeshie8):

(fog)(x) = f(g(x))

ganeshie8 (ganeshie8):

only that we need to remember, rest everything is donkey work :)

ganeshie8 (ganeshie8):

B. f(x)=5x^3, g(x)=x+1 (fog)(x) = f(g(x)) = f(x+1) = ? i want u complete it if u can

ganeshie8 (ganeshie8):

madam ...ttry ?

OpenStudy (anonymous):

sorry i was helping someone! I'm working on it right now @ganeshie8 :)

OpenStudy (anonymous):

@ganeshie8 so f(x)=5x and g(x)=(x+1)³ correct?

OpenStudy (anonymous):

nevermind, got the answer :)

ganeshie8 (ganeshie8):

hey sorry i slept off... glad u could get to the answer :)

OpenStudy (anonymous):

@ganeshie8 could you help me with the same answer please?

OpenStudy (anonymous):

because I dont think that (x)=5x and g(x)=(x+1)³ is the correct one